solve the system of equations
4x-4y=-8
y=2 + x

Answer by OHtutor(16)

4x=4y-8
y=2+x
x=(4y-8)/4=y-2
y=2+y-2=y
well y=y shows that y can be any number.
Are u sure that those are the corect equations?
Post the initial problem .